# The vertical structure of debris discs and the role of disc gravity

### v1

@@ 7 Summary
More generally, the current state of affairs in the modelling of planet–debris disk interactions warrants a cautious approach. Our results show that the disk’s gravity – often neglected without justification – plays an important dynamical role, even when the disk is less massive than the planet. At best, this oversight can compromise the accuracy of inferred planetary parameters; at worst, it leads to misleading interpretations of planetary system formation and evolution. A similar oversight persists in observational modelling, where Gaussian assumptions with constant $\mathcal{H}(R)$ are common for vertical profiles; we advocate instead for fitting non-Gaussian profiles and allowing for radial variations in aspect ratios. In closing, we note that this work lays the foundation for a more comprehensive numerical investigation, which we will undertake in future work, into the vertical structures imparted on non-zero-mass disks by planets.
@@ Acknowledgments
A.A.S. is supported by the Heising-Simons Foundation through a 51 Pegasi b Fellowship. The authors thank Joshua Lovell and Marija Jankovic for their engaging discussions, and Natasha Simonian for her assistance in enhancing the artistic quality of select figures. The results reported herein benefited from collaborations and/or information exchange within NASA’s Nexus for Exoplanet System Science (NExSS) research coordination network sponsored by NASA’s Science Mission Directorate and project “Alien Earths” funded under Agreement No. 80NSSC21K0593. K.M.K. acknowledges support under NASA Theoretical and Computational Astrophysical Networks (TCAN) via grant No. 80NSSC21K0497 and NASA XRP via grant No. 80NSSC24K0163. V.F.-G. acknowledges funding from the National Aeronautics and Space Administration through the Exoplanet Research Program under Grant No 80NSSC23K0288 (PI: V. Faramaz).
@@ Appendix A The Disk’s Back-reaction on the Planet: The Continuum Limit
and that the disk elements share a common longitude of ascending node with $\Omega_{d}(a_{d})={\rm cte}$. This approach and its assumptions are analogous to those in Sefilian et al. (2021), originally applied to coplanar setups (see their Appendix A).
@@
[…]
@@
[…]
@@
[…]
@@ Appendix B Damping of planetary inclinations due to (secular) resonant friction
[…]
@@
×$\qquad\qquad\times\big{[}b_{3/2}^{(1)}(\alpha_{\rm res})\big{]}^{2}\frac{n_{p}n_{d}(a_{\rm res})}{|dA_{p}/d\log a_{d}|_{a_{d}=a_{\rm res}}},$
@@
Finally, we point out that the damping rate given by Equation (B) matches that derived by Ward & Hahn (2003, see their equation (17)) using an alternative approach based on wave solutions in the tight-winding approximation. This is not surprising, as in the secular limit (i.e., when the pattern speeds are much smaller than the planetesimal’s mean motion), these results shall naturally follow from Toomre’s stellar dispersion relation (Toomre, 1969). Nevertheless, it is academically satisfying to see them derived from Lagrange’s planetary equations, as demonstrated here.
@@ D.1 Construction of Disk Surface Density Maps
To begin with, at a given time $t$, the orbit of each planetesimal (labeled $j=1,..,N$) is populated with $N_{np}$ new particles. Each new particle is assigned the same orbital elements as its parent planetesimal $j$ but with a mass $m_{j}/N_{np}$ and a mean anomaly $l$ randomly sampled from a uniform distribution between 00 and $2\pi$. This approach effectively imitates Gauss’ orbit-averaging technique (valid for secular dynamics; Tremaine, 2023), while simultaneously enhancing the resolution to $N_{t}\equiv N\times N_{np}\gg 1$ planetesimals.
@@ D.2 Scale Heights and Aspect Ratios
Here, the index $i$ counts the particles within a given radial bin centered at $R$, each with a vertical position $Z_{i}$ and mass $m_{i}$. Note that we compute $Z_{0}(R)$ since the vertical mass distribution of particles in the disk may not necessarily be symmetric around $Z=0$. However, we find that $Z_{0}(R)\approx 0$ in most cases and can be neglected – typically, $\mathrm{max}\big{(}|Z_{0}(R)|/R\big{)}\lesssim 0.5\%$ – except when resonant particles (if any) approach $R\rightarrow 0$, in which case $\mathrm{max}(|Z_{0}|)\sim 1$ au.
@@ D.3 Planet-disk misalignment
At a given time $t$, we assess the (mis)alignment between the planet and the disk using their angular momentum vectors. The angular momentum vector $\vec{L}_{k}$ of an object labeled $k$ (e.g., the planet $p$ or planetesimals $j=1,...,N$) reads as:
@@
$\theta_{{\rm al}}=\arccos\bigg{(}\frac{\vec{L}_{p}\cdot\vec{L}_{d}}{||\vec{L}_{p}||\cdot||\vec{L}_{d}||}\bigg{)},$ (76)

### v2

@@ 7 Summary
More generally, the current state of affairs in the modelling of planet–debris disk interactions warrants a cautious approach. Our results show that the disk’s gravity – often neglected without justification – plays an important dynamical role, even when the disk is less massive than the planet. At best, this oversight can compromise the accuracy of inferred planetary parameters; at worst, it leads to misleading interpretations of planetary system formation and evolution. A similar oversight persists in observational modelling, where Gaussian assumptions with constant $\mathcal{H}(R)$ are common for vertical profiles; we advocate instead for fitting non-Gaussian profiles and allowing for radial variations in aspect ratios.
In closing, we note that this work serves as a primer, offering a simplified analytical framework for exploring the role of disk gravity in misaligned planet–debris disk systems. Acknowledging its strengths and limitations, it provides a useful guide for identifying which planet–disc parameters and dynamical regimes merit the computational investment of more comprehensive studies, whether via N-body simulations or specialized secular “N-ring” approaches (e.g., Hahn, 2003; Touma et al., 2009; Sefilian et al., 2023).
@@ Acknowledgments
The authors thank Joshua Lovell and Marija Jankovic for their engaging discussions, Natasha Simonian for her assistance in enhancing the artistic quality of select figures, and the referee, Alexander Mustill, for a prompt report that improved the clarity of the manuscript. A.A.S. is supported by the Heising-Simons Foundation through a 51 Pegasi b Fellowship. The results reported herein benefited from collaborations and/or information exchange within NASA’s Nexus for Exoplanet System Science (NExSS) research coordination network sponsored by NASA’s Science Mission Directorate and project “Alien Earths” funded under Agreement No. 80NSSC21K0593. K.M.K. acknowledges support under NASA Theoretical and Computational Astrophysical Networks (TCAN) via grant No. 80NSSC21K0497 and NASA XRP via grant No. 80NSSC24K0163. V.F.-G. acknowledges funding from the National Aeronautics and Space Administration through the Exoplanet Research Program under Grant No 80NSSC23K0288 (PI: V. Faramaz).
@@ Appendix A The Disk’s Back-reaction on the Planet: The Continuum Limit
and that the disk elements share a common longitude of ascending node with $\Omega_{d}(a_{d})={\rm const}$. This approach and its assumptions are analogous to those in Sefilian et al. (2021), originally applied to coplanar setups (see their Appendix A).
@@
$\phi_{1}$ $=$ $\frac{1}{4}\bigg{(}\frac{a_{p}}{a_{\rm in}}\bigg{)}^{1-p}\int_{a_{p}/a_{\rm out}}^{a_{p}/a_{\rm in}}\alpha^{p-1}b_{3/2}^{(1)}(\alpha)d\alpha,$ (56)
@@
$\phi_{2}$ $=$ $\frac{1}{2}\bigg{(}\frac{a_{p}}{a_{\rm in}}\bigg{)}^{1-p-q}\int_{a_{p}/a_{\rm out}}^{a_{p}/a_{\rm in}}\alpha^{p+q-1}b_{3/2}^{(1)}(\alpha)d\alpha,$ (57)
@@
$\phi_{2}^{c}=\frac{1}{3}\frac{p+q+1}{1-\delta^{-p-q-1}}\frac{a_{\rm in}}{a_{p}}\int_{1}^{\delta}u^{-p-q-1}b_{3/2}^{(1)}\bigg{(}\frac{1}{u}\frac{a_{p}}{a_{\rm in}}\bigg{)}du.$ (58)
@@ Appendix B Damping of planetary inclinations due to (secular) resonant friction
$\frac{dI_{p}^{2}(t)}{dt}=-2\gamma I_{p}^{2}(t)=I_{p}^{2}(t)\frac{\pi}{4}n_{p}\frac{m_{p}}{M_{c}}\cdot{\rm Im}\bigg{[}\int_{L}F(a_{d})da_{d}\bigg{]}.$ (62)
@@
$\qquad\qquad\times\big{[}b_{3/2}^{(1)}(\alpha_{\rm res})\big{]}^{2}\frac{n_{p}n_{d}(a_{\rm res})}{|dA_{p}/d\log a_{d}|_{a_{d}=a_{\rm res}}},$
@@
In closing, we point out that the damping rate given by Equation (B) matches that derived by Ward & Hahn (2003, see their equation (17)) using an alternative approach based on wave solutions in the tight-winding approximation. This is not surprising, as in the secular limit (i.e., when the pattern speeds are much smaller than the planetesimal’s mean motion), these results shall naturally follow from Toomre’s stellar dispersion relation (Toomre, 1969). Nevertheless, it is academically satisfying to see them derived from Lagrange’s planetary equations, as demonstrated here. Finally, we note that secular resonant friction arises specifically from orbit-averaged torques and can be understood as a specialized form of the well-known dynamical friction process (Binney & Tremaine, 2008). Interestingly, a similar effect has been identified in the context of intermediate-mass black holes embedded in stellar disks around supermassive black holes (Szölgyén et al., 2021; Ginat et al., 2023).
@@ D.1 Construction of Disk Surface Density Maps
To begin with, at a given time $t$, the orbit of each planetesimal (labeled $j=1,..,N$) is populated with $N_{np}$ new particles. Each new particle is assigned the same orbital elements as its parent planetesimal $j$ but with a mass $m_{j}/N_{np}$ and a mean anomaly $l$ randomly sampled from a uniform distribution between 0 and $2\pi$. This approach effectively imitates Gauss’ orbit-averaging technique (valid for secular dynamics; Tremaine, 2023), while simultaneously enhancing the resolution to $N_{t}\equiv N\times N_{np}\gg 1$ planetesimals.
@@ D.2 Scale Heights and Aspect Ratios
Here, the index $i$ counts the particles within a given radial bin centered at $R$, each with a vertical position $Z_{i}$ and mass $m_{i}$. Note that we compute $Z_{0}(R)$ since the vertical mass distribution of particles in the disk may not necessarily be symmetric around $Z=0$. In practice, $Z_{0}(R)\approx 0$ in most cases, but we use Eq. (72) to maintain full generality; in particular it can account for offsets when resonant particles are excited to $R\rightarrow 0$.
@@ D.3 Planet-disk misalignment
At a given time $t$, we assess the (mis)alignment between the planet and the disk using their angular momentum vectors. The angular momentum vector $\vec{L}_{k}$ of an object labeled $k$ (e.g., the planet $p$ or planetesimals j=1,,Nj=1,...,N) reads as:
@@
[…]